OpenStudy (anonymous):

Help please! Determine whether the sequence is geometric. If it is, find the common ratio.

OpenStudy (anonymous):

\[1,\frac{ 1 }{ 2 },\frac{ 1 }{ 3 },\frac{ 1 }{ 4 }, ....\]